# GaMeS: Mesh-Based Adapting and Modification of Gaussian Splatting

### 02 Feb 2024 v1

@@ 4.2 Scenario I: Model with existing mesh
In this scenario, we utilize the provided mesh, incorporating Gaussians by strategically placing them onto its surface.
Our initial experiments used the Synthetic Dataset (Mildenhall et al., 2020), incorporating shared meshes. Table 1 presents PSNR metrics demonstrating the competitiveness of GaMeS approach with existing methods. Additional numerical comparisons are available in the supplementary material.
Notably, a significant contribution lies in the ease with which renders can be modified manually or in an automated process. Fig. 1 shows examples of reconstruction and simple object animations.
@@
4.3 Scenario II: Joint training of mesh and GS
Our approach is capable of modeling unbounded scenes. In order to achieve that, we first create an initial mesh and then adjust it during the training. Results in Table 2 demonstrate that such modeling is feasible and reasonably comparable to other existing models. Particularly, the results show parity with the GS-7K and R-SuGaR-15K models.
It is important to highlight that the rendering quality is contingent upon the quality of the acquired mesh. In our experiments, we opted for a hybrid approach to generating the initial mesh. This approach combines two meshes, each using a different number of Gaussian splats per face. The first mesh is extracted directly from the shared point cloud COLMAP, while the second mesh is inspired by the SuGaR technique (Guédon & Lepetit, 2023) and involves training GS over 7000 iterations to treat the 3D Gaussians positions as a point cloud. Only 20% of them are then transformed into a mesh using the Alpha Shapes (Edelsbrunner et al., 1983) algorithm.
As anticipated, we observe scene reconstruction, including both objects and details, see Fig. 2. Nevertheless, backgrounds and elements with large mesh faces may exhibit a lower-quality appearance. However, this does not hinder the feasibility of GaMeS model for reconstruction and modification. Additionally, the experiment revealed that Gaussian splats do not need to form a hierarchical structure. Instead, they can be embedded quite flatly, which is a non-trivial observation, and to our knowledge, it is not considered.
Numerous methods facilitate the generation of an initial mesh for objects, but our primary focus is not on this task. Instead, we emphasize addressing distinct issues that extend beyond the scope of initializing object meshes like EG3D (Chan et al., 2022), NGLOD (Takikawa et al., 2021b) or NeRFMeshing (Rakotosaona et al., 2023).
Figure 9: The main idea of fitting human mesh and how it can be transformed to a different person.
4.4 Scenario III: GaMeS with FLAME as Init Mesh
@@
Fig. 9, shows the main idea behind our approach. Initially, a human mesh is generated using the FLAME model. By accurately adjusting the mesh, we gain the ability to apply it, for example, to another person. By utilizing the FLAME parameters, we can modify various aspects, especially facial expressions, to generate facial expressions like smiling, sadness, eye closure, mouth opening, and flirty or disgusted expressions. The possibilities of reconstruction and modifications for all six faces are depicted in Fig. 8. All expressions were chosen randomly as our main goal was to show GaMeS model’s ability to render objects as well as its modifications/animations.
@@
The comparison of the mesh fitting capability between the GaMeS with the NeRFlame model is shown in Fig. 10. Notably, the faces are distinctly delineated and closely resemble the modeled objects, capturing not only postures but also expressions and intricate details.
Table 3 shows the results of using PSNR metrics on six faces. Comparisons are made between the two baseline NeRF models, GS, and the models created from them, i.e., the corresponding NeRFlame and GaMeS with animation capabilities.
@@
4.5 Real-time rendering
GaMeS GaMeS-A
FPS Num. of comp. PSNR FPS
bicycle 230 1 316 658 22.3 43.2
flowers 260 815 966 19.2 58.9
garden 163 2 092 692 25.9 31.9
stump 333 658 420 24.3 65.6
treehill 120 975 790 21.5 36.2
room 230 775 096 28.5 52.3
counter 213 502 808 25.9 60.0
kitchen 201 532 324 25.9 62.4
bonsai 174 1 274 648 27.0 42.4
Table 4: Comparison of achieved frame rate (FPS) on different scenes. GaMeS-A means FPS during modification. In each frame, we have different mesh positions. We can render static and dynamic scenes in real-time. All results were obtained using the RTX 4080 GPU.
We show that GaMeS can produce high-quality renders similar to GS or NeRF. Furthermore, we can manipulate mesh structures to obtain real-time modification of 3D objects, see Table 4. In our process of modification/simulation, we do not have any additional steps, which results in a considerable advantage of our approach.
To prove that, we conducted experiments to compare GS with GaMeS, examining their performance in static visualizations. We extend our experiments to include dynamic scenes by evaluating the performance of GaMeSin an animated scenario, see Fig. 2. By analyzing FPS metrics in static and animated contexts, we seek to provide real-time rendering. For comparability, all the computations were performed using the same RTX 4080 GPU.
@@ 5 Conclusion
This paper presents a new GS-based model called GaMeS that leverages two main concepts, i.e., classical mesh and vanilla GS. Here, we represent a 3D scene with a set of Gaussian components lying on the mesh. Moreover, we parameterize the Gaussian components by the mesh vertices, allowing single-stage training. In inference, we obtain GS, which can be modified in real time by manipulating mesh components. As a result, when the mesh changes, all Gaussian components automatically adapt to a new shape.
@@ Appendix A Appendix
Number of Gaussians
Chair Drums Lego Mic Materials Ship Hotdog Ficus
1
29.32 25.18 32.74 32.48 25.37 27.45 32.10 Z32.05
5
30.70 25.35 33.21 32.49 25.29 29.29 33.40 32.29
10
31.11 25.36 OOM 32.29 25.25 29.56 33.65 32.35
@@
1
0.066 0.052 0.027 0.012 0.060 0.133 0.062 0.018
5
0.050 0.049 0.022 0.012 0.061 0.107 0.042 0.018
10
0.044 0.049 OOM 0.12 0.063 0.102 0.038 0.019
Table 5: The quantitative comparisons (PSNR / SSIM / LPIPS) on Synthetic dataset with original meshes, with white background. OOM - CUDA out of memory. In this experiment we used Tesla V100-SXM2-32GB GPU.

### 15 Feb 2024 v3

@@ 4.2 Scenario I: Model with existing mesh
In this scenario, we utilize the provided mesh, incorporating Gaussians by strategically placing them onto its surface. We also allowed for moving vertices. Our initial experiments used the Synthetic Dataset (Mildenhall et al., 2020), incorporating shared meshes. Table 1 presents PSNR metrics demonstrating the competitiveness of GaMeS approach with existing methods. Additional numerical comparisons are available in the supplementary material. Notably, a significant contribution lies in the ease with which renders can be modified manually or in an automated process. Fig. 1 shows examples of reconstruction and simple object animations.
Figure 10: The main idea of fitting human mesh and how it can be transformed into a different person.
4.3 Scenario II: A model without mesh
When we do not have mesh, we train GS with flat Gaussians. Then, we construct the pseudo-mesh. Table 1 shows a comparison with other models on the Synthetic Dataset. Our approach is also capable of modeling unbounded scenes. In order to achieve that, we first create an initial pseudo-mesh and then adjust it during the training. Results in Table 2 demonstrate that such modeling is feasible and reasonably comparable to other existing models. Particularly, the results show parity with the GS-30K and R-SuGaR-15K models.
As anticipated, we observe scene reconstruction, including both objects and details, as can be seen in Fig. 2. Nevertheless, backgrounds and elements with large mesh faces may exhibit a lower-quality appearance. However, this does not hinder the feasibility of GaMeS model for reconstruction and modification.
4.4 Scenario III: GaMeS with FLAME as Init Mesh
@@
In Fig. 10, we show the main idea behind our approach. Initially, a human mesh is generated using the FLAME model. Thanks to accurately adjusting the mesh, we can apply it, for example, to another person. By utilizing the FLAME parameters, we can modify various aspects, especially facial expressions, to generate facial expressions like smiling, sadness, eye closure, mouth opening, and flirty or disgusted expressions. The possibilities of reconstruction and modifications for all six faces are depicted in Fig. 11. All expressions were chosen randomly as our main goal was to show GaMeS model’s ability to render objects as well as its modifications/animations.
@@
The comparison of the mesh fitting capability between the GaMeS with the NeRFlame model is shown in Fig. 12. Notably, the faces are distinctly delineated and closely resemble the modeled objects, capturing not only postures but also expressions and intricate details. Table 3 shows the results of using PSNR metrics on six faces. Comparisons are made between the two baseline NeRF models, GS, and the models created from them, i.e., the corresponding NeRFlame and GaMeS with animation capabilities.
@@ 5 Conclusion
This paper presents a new GS-based model called GaMeS that leverages two main concepts, i.e., classical mesh and vanilla GS. Here, we represent a 3D scene with a set of Gaussian components lying on the mesh. We can use the initial mesh as an input or estimate it during training. Moreover, we parameterize the Gaussian components by the mesh vertices, allowing single-stage training. In inference, we obtain GS, which can be modified in real-time by manipulating mesh components. As a result, when the mesh changes, all Gaussian components automatically adapt to a new shape.
@@ Appendix A Appendix
Number of Gaussians Chair Drums Lego Mic Materials Ship Hotdog Ficus
1 29.32 25.18 32.74 32.48 25.37 27.45 32.10 Z32.05
5 30.70 25.35 33.21 32.49 25.29 29.29 33.40 32.29
10 31.11 25.36 OOM 32.29 25.25 29.56 33.65 32.35
pseudo-mesh 35.34 25.93 35.42 35.34 30.01 29.58 37.65 34.81
@@
1 0.066 0.052 0.027 0.012 0.060 0.133 0.062 0.018
5 0.050 0.049 0.022 0.012 0.061 0.107 0.042 0.018
10 0.044 0.049 OOM 0.012 0.063 0.102 0.038 0.019
pseudo-mesh 0.012 0.039 0.015 0.006 0.033 0.105 0.019 0.012
Table 4: The quantitative comparisons (PSNR / SSIM / LPIPS) on Synthetic dataset with original meshes or pseudo-mesh, with white background. OOM - CUDA out of memory. In this experiment we used Tesla V100-SXM2-32GB GPU.
@@ A.2 Modification
Fig. 17 shows that modifications are possible even using only pseudo-mesh.
@@
Figure 17: An example of modification using the GaMeS model with pseudo-mesh.